# SegAnyPET: Universal Promptable Segmentation from Positron Emission Tomography Images

### v3

@@ 4.2 Experimental Results
Comparison with Universal Foundation Models. The original SAM [15] and MedSAM [21] are designed for 2D segmentation tasks and cannot handle 3D inputs directly, necessitating input prompts for each 2D slice containing the target to conduct slice-by-slice segmentation. In contrast, SegAnyPET and other 3D SAM medical adaptions [34] can be directly utilized to segment the target organs from input volume with one or few prompts. Table 1 presents the performance of SegAnyPET with state-of-the-art foundation models on our internal test set under different prompt settings. We can observe that 2D models exhibit lower performance due to the lack of volumetric spatial information. In contrast, 3D models can achieve better performance with less manual prompting effort. However, all these existing foundation models show poor generalization performance on PET segmentation task. Compared with state-of-the-art methods, SegAnyPET exhibits significantly better segmentation performance and outperforms the second-ranked method by a significant improvement up to 18.40% on average DSC. Besides, we visualize the segmentation results of SegAnyPET and other segmentation foundation models for organ segmentation in Fig. 4, which intuitively show that SegAnyPET performs more precise segmentation with higher DSC using the same or fewer manual efforts.
@@
Comparison with Task-Specific Models. In addition to foundation models, we also compare with state-of-the-art task-specific models in Tab. 1. These task-specific models are trained for the segmentation of target organs. When adapting to training invisible organs, additional annotation and training are required. We conduct two experimental settings, including training on HQ data alone and training with the combination of HQ and LQ data. Our observations show that incorporating additional LQ data does not consistently enhance model performance and may even degrade the performance in some scenarios. Despite the efforts of prompting, SegAnyPET demonstrates superior performance while retaining generalization capability to unseen targets.
@@
Generalization to Unseen Organs. To validate the generalization performance of SegAnyPET, we conduct additional experiments of foundation models for promptable segmentation of training invisible organs. As shown in Tab. 2, we observe SegAnyPET also exhibits strong generalization performance and outperforms state-of-the-art segmentation foundation models. The experimental results demonstrate that by efficiently learning representations from large-scale PET images, SegAnyPET exhibits comparable performance for training invisible organs. We further evaluate SenAnyPET on training invisible AutoPET and AutoPET-Organ dataset with significant distribution gap compared with PETS-5k dataset. The results and analysis are shown in the Appendix.
@@
Ablation Analysis. In Tab. 3, we conduct ablation experiments to evaluate the effectiveness of proposed training strategy. Three different training strategies are conducted including fine-tuning with mixed HQ and LQ, training with additional unsupervised consistency regularization, and the proposed CPCL strategy. Experimental results show that the proposed strategy can achieve better performance on both seen and unseen targets.
5 Discussion & Conclusion
@@
In this work, we propose SegAnyPET, a modality-specific foundation model for universal promptable segmentation from 3D PET images. To this end, we collect and construct a large-scale PET segmentation dataset, PETS-5k, which consists of 5,731 three-dimensional whole-body PET images, encompassing over 1.3M 2D images to train the foundation model. To facilitate the model training with different annotation quality, we design a cross prompting confident learning strategy building on the foundation of promptable segmentation for noise-robust learning. Our model shows substantial capabilities in segmenting both training visible and invisible target organs using only one or few prompt points. Its performance significantly exceeds both state-of-the-art segmentation foundation models and task-specific fully supervised models.
While SegAnyPET boasts strong capabilities, there are still inherent limitations due to manual efforts for promptable segmentation, which could potentially reduce the usability of the model compared with fully automatic segmentation. We aim to incorporate semantic information together with positional prompts as additional support to enable efficient and precise auto-prompting for automatic segmentation in the future. We also aim to enlarge existing data and task to more comprehensive scenarios including different tracers and tumor lesion segmentation. In conclusion, as the first foundation model for PET images, we believe that SegAnyPET will advance the developments of universal segmentation and serve as a powerful pre-trained model to facilitate the applications to various downstream tasks for molecular imaging.
@@
Table 3: Ablation analysis of different training strategies for interactive segmentation from PET images.
Acknowledgement
This work was supported by the National Natural Science Foundation of China (Grant No. 82394432 and 92249302), and the Shanghai Municipal Science and Technology Major Project (Grant No. 2023SHZDZX02). The computations in this research were performed using the CFFF platform of Fudan University.
@@ A.1. PETS-5k Dataset
The proposed PETS-5k dataset consists of 5,731 three-dimensional whole-body 18F-FDG PET images collected from one local medical center. Patients were fasted for at least 6h and had a blood glucose level $<$ 200 mg/dL before the PET/CT examination. PET/CT imaging was performed at a median uptake time of 67 min (range from 53 to 81 min) after intra-venous injection of 18F-FDG (3.7 MBq/kg). All data were acquired on PET/CT scanners (Siemens Biograph mCT) with 5 min per bed position. A low-dose CT scan (120 kVp; 40–100 mAs; 5 mm slice thickness) was performed from the upper thigh to the skull base, followed by a PET scan with 3D Flowmotion acquisition mode. PET images were reconstructed with 4.07 $\times$ 4.07 $\times$ 3 $mm^{3}$ voxels using CT-based attenuation correction by Siemens-specific TrueX algorithm.
@@ A.4. Summary and Visualization
The information summary and visualization of datasets used in our work are shown in Tab. 4 and Fig. 6.
@@ B.1. SegAnyPET Architecture
The detailed architecture of SegAnyPET is shown in Fig. 7. Following the design in [34], for the image encoder, the input patch size is set to 16$\times$16$\times$16 with a patch embedding dimension of 768, paired with a learnable 3D absolute positional encoding. Then the embeddings of patches are input to 3D self-attention blocks. The depth of self-attention blocks is set to 16. Within the prompt encoder, sparse prompts are leverage by 3D position embedding to represent 3D spatial differences, while dense prompts are handled with 3D convolutions followed by layer normalization and GELU activation. The mask decoder is integrated with 3D upscaling procedures, employing 3D transformer blocks and 3D transposed convolutions to get the final segmentation result.
@@ B.2. Implementation Training Details
Our method is implemented in Python with PyTorch and trained on 4 NVIDIA Tesla A100 GPUs, each with 80GB memory. We use the AdamW optimizer with an initial learning rate of 0.0008 and a weight decay factor of 0.1. The training was performed for a total of 200 epochs on the constructed PETS-5k dataset. The batch size is set to 12 with a volumetric input patch size of 128$\times$128$\times$128. To handle the learning rate schedule, we employed the MultiStepLR scheduler, which adjusts the learning rate in predefined steps with 120 and 180 epochs, with a gamma value of 0.1, indicating that the learning rate is reduced by 10% of its original value at each step. In distributed training scenarios, we utilized gradient accumulation with 20 steps to simulate larger effective batch sizes, which can improve model performance by providing a more accurate estimate of the gradient. For total loss in the training loop, the ramp-up trade-off weighting coefficient $\lambda$ is scheduled by the time-dependent Gaussian function as $\lambda=\omega_{max}*e^{-5(1-t/t_{max})}$, where $t_{max}$ is the maximum training iteration, $\omega_{max}$ is the maximum weight set as 0.1 and $\beta$ is set to 5. The weighting coefficient can avoid the domination by misleading targets at the early training stage.
@@ B.4. Evaluation Metric
where $G$ represents the ground truth segmentation and $S$ represents the predicted segmentation. DSC ranges from 0 to 1, where 1 indicates perfect overlap between the ground truth and the predicted segmentation.
Method SAM [15] MedSAM [21] SAM-Med3D [34] SegAnyPET
Lung-LL 13.16 15.49 21.93 2.84 21.81 22.19 32.27 38.05 41.77 13.32 24.09 26.73
Table 5: Generalization performance to unseen out-of-distribution AutoPET-Organ dataset with comparison to state-of-the-art general-propose foundation models for interactive segmentation from PET images.
Modality Registrated CT Registrated CT PET PET
Avg DSC 88.71 66.59 73.77 76.01 72.09 76.58 78.35 90.49 90.90 91.05
Table 6: Quantitative comparison different automatic and promptable segmentation models for organ segmentation from CT and PET images.
@@ C. Additional Experiments and Discussion
C.1. Comparison with Segmentation from CT
Since the data used in this work are whole-body PET/CT images, we conduct additional evaluations with models for automatic and promptable organ segmentation from CT images. Given that the segmentation is used to evaluate the organ metabolic intensity from PET, it is necessary to register the CT to the resolution of PET before segmentation. We compare SegAnyPET with state-of-the-art CT segmentation model TotalSegmentator [36] and SAM-Med3D-turbo applied to both registered CT and PET images Through the experimental comparison between the registered CT and PET in Tab. 6, we observe that the performance of segmentation from CT is inferior to that of direct segmentation from PET. This performance gap, combined with the long-standing and significant concern of reducing radiation exposure, underscores the importance of developing a PET-only segmentation model which could be compatible to diverse scenarios, including PET/MRI or CT-free PET with self-attenuation correction.
As PET images often come with corresponding structural CT and MRI images for attenuation correction, we plan to extend SegAnyPET to a multi-modal scenario by utilizing the information from these modalities to assist in the segmentation procedure, thereby improving the overall performance of the segmentation process.
C.2. Generalization to Unseen Dataset for Organ Segmentation
We evaluated the generalization ability of SegAnyPET to unseen out-of-distribution AutoPET-Organ dataset with comparison to state-of-the-art general-propose foundation models for interactive segmentation. As shown in Tab. 5, we observe that SegAnyPET achieves satisfying generalization ability on OOD data and outperforms existing models across different prompt settings for training-visible organs. For training-invisible organs, while it surpasses SAM and MedSAM, it lags behind SAM-Med3D. One possible explanation is that SegAnyPET is trained on relatively few targets with 5 categories. In contrast, SAM-Med3D is trained on a large amount of 245 categories including multiple organs and lesions. Some unseen targets for SegAnyPET like aorta, lung lobes and prostate are visible to SAM-Med3D from other modalities. We observe when generalized to unseen targets or training-visible targets on OOD data, SegAnyPET outperforms existing models. However, when encountering both data and target shift, SegAnyPET exhibits slightly weaker performance. We aim to improve current design with techniques like using SuperVoxel to enhance training label diversity [10] in future work.
C.3. Preliminary Analysis on Tumor Segmentation
In addition to the internal and external evaluation on organ segmentation, we also conduct a preliminary evaluation SegAnyPET with other state-of-the-art segmentation foundation models for zero-shot tumor segmentation on AutoPET dataset. Table 7 presents the experimental results under different prompt settings. Contrary to the conclusions of organ segmentation, we observe that slice-by-slice segmentation of 2D foundation models outperforms 3D foundation models. A significant difference is that the target organs in our task are all continuous entities, while the whole-body tumors in AutoPET dataset are scattered multiple small targets located in various places, as shown in Fig. 9. Therefore, the 3D model cannot directly segment all these scattered tumors using one or a few prompt points. As an important application scenario, we aim to enlarge the dataset with instance-level tumor annotation for training and evaluation of tumor lesion segmentation in the future.

### v4

@@ 4.2 Experiments
Comparison with Universal Foundation Models. The original SAM [15] and MedSAM [21] are designed for 2D segmentation tasks and cannot handle 3D inputs directly, necessitating input prompts for each 2D slice containing the target to conduct slice-by-slice segmentation. In contrast, SegAnyPET and other 3D SAM medical adaptions [34] can be directly utilized to segment the target organs from input volume with one or few prompts. Table 1 presents the performance of SegAnyPET with state-of-the-art foundation models on our internal test set under different prompt settings. We can observe that 2D models exhibit lower performance due to the lack of volumetric spatial information. In contrast, 3D models can achieve better performance with less manual prompting effort. Compared with state-of-the-art methods, SegAnyPET exhibits significantly better segmentation performance and outperforms the second-ranked method by a significant improvement up to 18.40% on average DSC. We visualize the segmentation results of SegAnyPET and other segmentation foundation models for organ segmentation in Fig. 4, which intuitively show that SegAnyPET performs more precise segmentation with higher DSC.
Comparison with Task-Specific Models. In addition to foundation models, we also conduct a comparison with state-of-the-art task-specific models in Tab. 1. These task-specific models are trained on the HQ training set focusing on the segmentation task of target organs. When adapting to training invisible organs, additional annotation and model training are required for segmentation. Despite the manual prompting efforts, we observe that SegAnyPET significantly outperforms task-specific models, while preserving the generalization ability for universal segmentation.
@@
Method SAM [15] MedSAM [21] SAM-Med3D [34] SegAnyPET
Lung-$L$ 13.16 15.49 21.93 2.84 21.81 22.19 32.27 38.05 41.77 13.32 24.09 26.73
Table 3: Generalization performance to unseen out-of-distribution AutoPET-Organ dataset with comparison to state-of-the-art segmentation foundation models for zero-shot interactive segmentation from PET images.
@@
Generalization to Unseen Organs and Dataset. To validate the generalization performance of SegAnyPET, we conduct additional experiments of foundation models for promptable segmentation of training invisible organs. As shown in Tab. 2, we observe SegAnyPET also exhibits strong generalization performance and outperforms state-of-the-art segmentation foundation models. The experimental results demonstrate that by efficiently learning representations from large-scale PET images, SegAnyPET exhibits comparable performance for training invisible organs. Furthermore, we evaluate the proposed method on training invisible AutoPET-Organ dataset with significant distribution gap compared with PETS-5k dataset. As shown in Tab. 3, we observe that SegAnyPET can achieve satisfying generalization ability on unseen dataset.
@@
Ablation Analysis. In Tab. 4, we conduct ablation experiments to evaluate the effectiveness of proposed training strategy. Three different training strategies are conducted including fine-tuning with mixed HQ and LQ, training with additional unsupervised regularization for LQ, and the proposed cross prompting confident learning strategy. Experimental results show that the proposed strategy can achieve better performance on both seen and unseen targets.
5 Discussion and Conclusion
@@
In this work, we propose SegAnyPET, a modality-specific foundation model for universal promptable segmentation from 3D PET images. To this end, we collect and construct a large-scale PET segmentation dataset, PETS-5k, which consists of 5,731 three-dimensional whole-body PET images, encompassing over 1.3M 2D images to train the foundation model. To facilitate the model training with different annotation quality, we design a cross prompting confident learning strategy building on the foundation of promptable segmentation for noise-robust learning. Through comprehensive evaluations on both internal and external datasets, our model shows substantial capabilities in segmenting both seen and unseen targets using only one or few prompt points and robust generalization abilities to manage new data and tasks. Its performance significantly exceeds both state-of-the-art segmentation foundation models and task-specific fully supervised models.
@@
Table 4: Ablation analysis of different training strategies for zero-shot interactive segmentation from PET images.
While SegAnyPET boasts strong capabilities, there are still inherent limitations due to manual efforts for promptable segmentation, which could potentially reduce the usability of the model compared with fully automatic segmentation. We aim to incorporate semantic information together with positional prompts as additional support to enable efficient and precise auto-prompting for automatic segmentation in the future. Besides, despite the differences between existing structural medical images, these large-scale medical datasets [1] can still provide useful guidance for training in conjunction with PET images. In conclusion, as the first foundation model for PET images, we believe that SegAnyPET will advance the developments of universal segmentation and serve as a powerful pre-trained model to facilitate the applications to various downstream tasks for molecular imaging.
@@ A.1. PETS-5k Dataset
The proposed PETS-5k dataset consists of 5,731 three-dimensional whole-body 18F-FDG PET images collected from one local medical center. Patients were fasted for at least 6h and had a blood glucose level << 200 mg/dL before the PET/CT examination. PET/CT imaging was performed at a median uptake time of 67 min (range from 53 to 81 min) after intra-venous injection of 18F-FDG (3.7 MBq/kg). All data were acquired on PET/CT scanners (Siemens Biograph mCT) with 5 min per bed position. A low-dose CT scan (120 kVp; 40–100 mAs; 5 mm slice thickness) was performed from the upper thigh to the skull base, followed by a PET scan with 3D Flowmotion acquisition mode. PET images were reconstructed with 4.07 $\times$ 4.07 $\times$ 3 $mm^{3}$ voxels using CT-based attenuation correction by Siemens-specific TrueX algorithm.
@@ A.4. Summary and Visualization
An overview information summary and visualization of the datasets used in our work are shown in Tab. 5 and Fig. 5.
@@ B.1. SegAnyPET Architecture
The detailed architecture of SegAnyPET is shown in Fig. 6. Following the design in [34], for the image encoder, the input patch size is set to $\times16$$\times$16 with a patch embedding dimension of 768, paired with a learnable 3D absolute positional encoding. Then the embeddings of patches are input to 3D self-attention blocks. The depth of self-attention blocks is set to 16. Within the prompt encoder, sparse prompts are leverage by 3D position embedding to represent 3D spatial differences, while dense prompts are handled with 3D convolutions followed by layer normalization and GELU activation. The mask decoder is integrated with 3D upscaling procedures, employing 3D transformer blocks and 3D transposed convolutions to get the final segmentation result.
@@ B.2. Implementation Training Details
Our method is implemented in Python with PyTorch and trained on 4 NVIDIA Tesla A100 GPUs, each with 80GB memory. We use the AdamW optimizer with an initial learning rate of 0.0008 and a weight decay factor of 0.1. The training was performed for a total of 200 epochs on the constructed PETS-5k dataset. The batch size is set to 12 with a volumetric input patch size of $\times128$$\times$128. To handle the learning rate schedule, we employed the MultiStepLR scheduler, which adjusts the learning rate in predefined steps with 120 and 180 epochs, with a gamma value of 0.1, indicating that the learning rate is reduced by 10% of its original value at each step. In distributed training scenarios, we utilized gradient accumulation with 20 steps to simulate larger effective batch sizes, which can improve model performance by providing a more accurate estimate of the gradient. For total loss in the training loop, the ramp-up trade-off weighting coefficient $\lambda$ is scheduled by the time-dependent Gaussian function as $\lambda=\omega_{max}*e^{-5(1-t/t_{max})}$, where $t_{max}$ is the maximum training iteration, $\omega_{max}$ is the maximum weight set as 0.1 and $\beta$ is set to 5. The weighting coefficient can avoid the domination by misleading targets at the early training stage.
@@ B.4. Evaluation Metric
where $G$ represents the ground truth segmentation and $S$ represents the predicted segmentation. The DSC is calculated by taking twice the size of the intersection and dividing it by the sum of the sizes of the two sets. This normalization ensures that the coefficient ranges from 0 to 1, where 1 indicates perfect overlap between the ground truth and the predicted segmentation.
@@ C. Additional Experiments and Discussion
C.1. Generalization to Tumor Segmentation
@@
In addition to the internal and external evaluation on organ segmentation, we also compare SegAnyPET with other state-of-the-art segmentation foundation models for zero-shot tumor segmentation on AutoPET dataset. Table 6 presents the experimental results under different prompt settings. Contrary to the conclusions of organ segmentation, we observe that slice-by-slice segmentation of 2D foundation models outperforms 3D foundation models. A significant difference is that the target organs in our task are all continuous entities, while the whole-body tumors in AutoPET dataset are scattered multiple small targets located in various places, as shown in Fig. 8. Therefore, the 3D model cannot directly segment all these scattered tumors out using one or a few prompt points, which also indicates the limitations of our method for this task. As an important application scenario, we aim to enlarge the dataset with instance-level tumor annotation for training and evaluation of tumor lesion segmentation in the future.
Modality Registrated CT Registrated CT PET PET
Avg DSC 88.71 66.59 73.77 76.01 72.09 76.58 78.35 90.49 90.90 91.05
Table 7: Quantitative comparison different automatic and promptablt segmentation models for organ segmentation from CT and PET images.
C.2. Comparison with CT Segmentation
Since the data used in this work are whole-body PET/CT images, we conduct additional evaluations with models for automatic and promptable organ segmentation from CT images. Given that the segmentation is used to evaluate the organ metabolic intensity from PET, it is necessary to register the CT to the resolution of PET before segmentation. Through the experimental comparison between the registered CT and PET in Tab. 7, we observe the performance of segmentation from CT is inferior to that of direct segmentation from PET. Furthermore, developing a PET segmentation model can be compatible to different scenarios such as PET/MRI or CT-free PET with self-attenuation correction.